OpenStudy (anonymous):

Write three fractions that are equivalent to 6/9 . Show your work.

OpenStudy (anonymous):

6/9 x 2/2 = 12/18 there's one 6/9 x 3/3 = 18/27 there's another 6/9 x 4/4 = 24/36 last one

OpenStudy (anonymous):

thanks!